• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name Chao

    Log entry time 16:00:46 on May 03, 2012

    Entry number 381930

    keyword=Day shift summary

    Shift workers:
    Chao Gu(SL), Navaphon Muangma(TO), Patricia Solvignon-Slifer(3rd)

    08:00 Shift starts. Pengjia is doing slow raster calibration.
    10:30 Pengjia finished his beam line calibration. We start to center beam.
    11:30 Right septa tripped at 900A. Jack came here and reset it. The new limit for right septa is 850A. We are not able to do optics at this setting.
    12:30 Right septa tripped again at 840A. Jack reset it again and set the new limit to be 800A. Meanwhile MCC is doing some trouble shooting work on C100 cavities.
    13:15 We move LHRS to 2.2510GeV (Septa 725.4A) RHRS to 1.7348GeV (Septa 724.6A) for safe.
    13:30 No beam in the Hall. MCC is checking some IOC problem.
    15:30 Escorted access for multiple trouble shooting work. Heidi is working on left arm's dipole valve. Jack is trying to find out the reason of the septa trip.
    16:00 Shift ends. We still do not have beam in the hall. Heidi and Jack are still in the hall to do their work.